Algebraic multilevel methods for porous media flow problems

Jürgen Fuhrmann WIAS Berlin, Germany

Abstract

An algebraic method for the construction of multilevel preconditioners is discussed which puts algebraic multigrid methods on both structured and unstructured meshes into one framework. Based on implementations in both the structured and the unstructured cases, the efficiency of the method is tested. Furthermore, the performance of the method in an adaptive context is discussed.

This method is applied within a framework for the numerical solution of nonlinear parabolic equations, e.g. Richard’s equation or the equation describing the movement of a viscous compressible fluid in a porous medium.